OpenStudy (s):

Square root of (2/3)^2( square root(2/3)^2-4)= ?

OpenStudy (mertsj):

\[\sqrt{(\frac{2}{3})^2}\sqrt{(\frac{2}{3})^2}-4\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (mertsj):

It is much easier to help someone who will actually respond to one's questions.